### v1

# Winning opinion: Following Your Friends’ Advice or That of Their Friends?

Abstract
We investigate a variation of the classical voter model in which the set of influencing agents depends on an individual’s current opinion. The initial population consists of a random sample of equally sized subpopulations for each state, and two types of interactions are considered: (i) direct neighbors, and (ii) second neighbors (friends of direct neighbors, excluding the direct neighbors themselves). The neighborhood size, reflecting regular network connectivity, is kept constant across all agents.
Our findings reveal that varying the interaction range introduces asymmetries that influence the probability of consensus and convergence time. At low connectivity, direct neighbor interactions dominate, driving consensus. As connectivity increases, the probability of consensus for either state becomes equal, mirroring symmetric dynamics. Asymmetry is also reflected in convergence time: while symmetric cases show decreasing times with increased connectivity, asymmetric cases exhibit an almost linear increase. The introduction of stubborn agents further accentuates these effects, particularly when they favor the less dominant state, significantly increasing consensus time.
We conclude by discussing the implications of these findings for decision-making processes and political campaigns in human populations.
I Introduction
Human behavior, especially the formation and evolution of opinions, is influenced by a complex interplay of factors, many of which remain poorly understood [1, 2]. Among these factors, social ties play a crucial role in shaping individual opinions and preferences.
Humans inherently form networks through which information is exchanged. These networks can vary significantly depending on the context. For instance, connections formed among university peers may differ from those formed in romantic or social contexts. Nonetheless, individuals often participate in multiple overlapping networks [3, 4].
The way an individual’s “neighborhood of influence” is defined can significantly affect opinion dynamics. Even when network structures are uniform, the interaction rules can determine which opinions prevail (see, for example, [5, 6, 7]). This study explores whether individuals benefit more by following the opinions of their immediate friends or their friends’ friends.
The question was inspired by a personal observation: “When I evaluate the potential success of a publication, I find that my friends’ feedback often lacks objectivity. In contrast, their friends’ assessments, though less personal, tend to be more accurate and unbiased.” Motivated by this, we investigate the dynamics of opinion formation under two scenarios: when agents are influenced by direct friends or by friends of friends.
In this study, we simulate opinion dynamics in a population of $N$ agents on a regular network with circular boundary conditions. Each agent can adopt one of two states (e.g., 0 or 1). Agents in state 0 are influenced by their direct neighbors, while those in state 1 are influenced by their second neighbors (excluding direct neighbors). Using the classical voter model [8, 9, 10], with constant size and asynchronous updates, we analyze the probability of consensus and the time required to achieve it under symmetric and asymmetric neighborhood definitions.
Figure 1: Opinion dynamics occur on a regular network as depicted in this figure. The central inset shows the different neighborhoods that are considered in this model: (Top) Nearest neighbors and (Bottom) Second neighbors (excluding the common neighbors). In this example, the network size is: $N=64$ and the connectivity is: $k=4$ (for both neighborhoods). The maximum connectivity that assures no intersection between both neighborhoods is $k_{max}=16$.
II Voter model on a regular cyclic network
We consider a population of $N$ agents, each adopting one of two states (0 or 1), arranged on a regular cyclic network with connectivity $k$. Initially, the population is evenly divided between states 0 and 1, with agents randomly distributed across the network. Simulations are carried out until a final time of 1 million, guaranteeing that the population reaches one of the two consensus state, either 0 or 1. Note that the maximum connectivity that assures that the two neighborhoods First-First and First-Second have no common agents is $k_{max}=\frac{1}{2}(N-2)$.
According to the classical Voter model [11], at each time step, a node is randomly selected, and its state is compared to that of another randomly chosen node within its neighborhood. If the two states differ, the first node adopts the second node’s state (see Fig. 1). One time step is defined as $N$ updates, ensuring that, on average, each node is updated once.
In the symmetric case, neighborhoods are identical for both states ($k$-nearest neighbors). In the asymmetric case, neighborhoods depend on the agent’s state: state 0 agents interact with nearest neighbors, while state 1 agents interact with their second neighbors (avoiding the nearest neighbors of the central node). We analyze the probability of reaching a consensus (all agents in one state) and the time required for this consensus under both scenarios.
In the following subsections, we present results obtained about: (i) the probability of reaching a given consensus, either 0 or 1 and (ii) the time the population takes to achieve such a consensus. The last subsection shows the results obtained when one of the agents keep its state forever, a so called stubborn or zealot, with regards to (i) and (ii) measures.
II.1 Probability of reaching a consensus population
[…]
This probability of reaching the 0-state consensus depends on the network size $N$ Fig. 2b. However, when plotted against normalized connectivity $\frac{k}{N-1}$, these probabilities collapse onto a single curve across different population sizes, indicating a universal dependence on the normalized parameter, as shown in Fig. 2c.
Figure 2: (a). Probability of achieving a consensus population of individuals at state 0, $P(X=0)$, as a function of the network connectivity $k$ for the two cases: (blue) symmetric population, i.e. both states have the same nearest neighborhood (First-First), and (red) asymmetric population where individuales at state 0 interact with their nearest neighbors, while individuals at state 1 interact with their second neighbors (avoiding first neighbors). In contrast to the symmetric case, where this probability is independent of $k$, when each state has a different neighborhood, $P(X=0)=1$ for low values of the connectivity and decreases monotonously to 1/2 as $k$ increases, approaching the symmetric dynamics. The probability is estimated from 100 simulations, and the network size is $N=200$. (b) The same probability $P(X=0)$ in terms of $k$ for three networks sizes (estimated from the same number of simulations) The maximum connectivity that assures that the two neighborhoods do not intercept is $k_{max}=\frac{1}{2}(N-2)$. (c) As the above pannel but normalizing the $X$-axis by the maximun connectivity $N-1$. As seen, the three curves collapse in one, showing the asymptotic decrease to the symmetric probability $P(X=0)=1/2$ as $k$ increases.
II.2 Consensus time
Figure 3 shows the results obtained from the simulations described in the previous section, where the consensus time for each state was recorded. Consensus time is defined as the first step at which the entire population adopts a single state (either 0 or 1), and it is calculated separately for each state. The reported consensus times represent the average of all simulations that result in a particular state.
For large connectivity values, the consensus times for both states are similar. However, for low connectivity values in the asymmetric case, consensus is only achieved for state 0. As shown in Figure 3a, for the symmetric case (where both states have identical neighborhoods), the consensus time decreases with increasing network connectivity, following a quasi-exponential trend. In contrast, for the asymmetric case, the consensus time increases with connectivity, as illustrated in Figure 3b.
Figure 3: Time steps to a consensus population as a function of the network connectivity $k$. Asynchronous updating on a network formed by 200 nodes (agents). (a) Symmetric case: both states have the same nearest neighborhood, (b) Asymmetric case: nodes at state 0 interact with their nearest neighbors and nodes at state 1 interact with their second neighbors, avoiding the common neighbors. Blue curve corresponds to the 0-state consensus and red curve to 1-state consensus. Note that for low $k$-values only red points exist since reaching the 0-state consensus has probability 1. Importantly, the dependence of the consensus time with $k$ differs drastically between the two figures, while it descreases (almost exponentially) for the symmetric case, it increases (almost linearly) with $k$ for the asymmetric case. Moreover, as it can be observed, the consensus time in the asymmetric case is lower than in the symmetric one.
The behavior shown in Fig. 3b for the asymmetric case, with a network size of $N=200$, is consistent across other network sizes (see Fig. 4). For low $k$ values, where consensus exclusively occurs in the 0-state population, the functions representing the time to consensus (measured as time steps divided by network size) increase linearly with the normalized network connectivity, $\frac{k}{N-1}$. Notably, the slope of this linear increase becomes steeper for larger network sizes, $N$.
Figure 4: Time until consensus, i.e. time steps divided by $N$, for an initial population to reach a consensus, as a function of the network connectivity $k$ for the asymmetric cases for different network sizes: 200 (blue), 100 (red) and 50 (green). Light colors correspond to the time to 1-state consensus and dark colors to 0-state consensus. Note that for low $k$-values for the three $N$ sizes, only dark points appear since only consensus around the 0-state occurs. For this state, as it can be observed, the consensus time exhibits a minimun value for the three sizes, at a connectivity (normalized) value that depends on $N$. The three curves can be fitted to a straight line: $tau=m\,k+n$, with the same slope $m$ and the intercept $n=f(N)$. The maximum connectivity that assures that the two neighborhoods do not intercept is $k_{max}=\frac{1}{2}(N-2)$.
The discrepancy between the symmetric and asymmetric cases regarding how consensus time depends on connectivity is not entirely understood. However, it can be attributed to competition between the different neighborhood structures, which creates an asymmetry in the probabilities of reaching consensus. Specifically, the probability of achieving consensus in state 0 is higher than that for state 1 (see Fig. 2a). This asymmetry in probabilities results in a shorter time to reach consensus, particularly at low $k$-values.
As connectivity ($k$) increases, the probabilities of reaching consensus for states 0 and 1 converge, as does the time required to achieve consensus (Fig. 3b). It is also noteworthy that for $k>10$, the consensus times for both states are effectively equal (see the red and blue curves in Fig. 3b). For $k<10$, where the probability of reaching consensus in state 0 is nearly 1, the consensus time increases. This behavior is similar to the symmetric case, where both states have identical neighborhoods.
Figure 5: Time to consensus for both states in the presence of a stubborn agent (node), either in state 0 (blue) or 1 (red) in a network with $N=100$ nodes. Note that for $k=2,4,6$ the consensus time when the state of the stubborn agent is 1 is not depicted due to its extremely large value (larger than $10^{7}$ time steps). As in the asymmetric case without stubborn agents, the time to 1-consensus is larger than the corresponding time for the 0-consensus for all $k$-values, although these times converge as $k$ increases.
III Effects of stubborn agents
To better understand the effect of neighborhood asymmetry on the opinion dynamics of the population, we consider the inclusion of agents whose opinions remain fixed over time, known as stubborn agents. It is well established that the presence of stubborn agents supporting both states can prevent the population from achieving consensus [13].
In the simplest case of a single stubborn agent, the dynamics of the classical voter model with nearest neighbors are biased toward the agent’s state. For instance, if the stubborn agent permanently holds the 0-state, the entire population will eventually converge to this state, regardless of its initial configuration [14]. In essence, the stubborn agent’s state dominates and overrides the other state. The time required to achieve this consensus can also be determined in the infinite population limit [14].
In asymmetric neighborhoods, the effect of a stubborn agent remains similar to that in the symmetric case: the agent drives the population to consensus in its own state. However, the time required to reach this consensus is significantly affected by the presence of the stubborn agent (see Fig. 5). If the stubborn agent holds the 0-state (nearest neighbor interactions), the consensus time is significantly reduced for any value of $k$. Conversely, if the stubborn agent holds the 1-state (second-neighbor interactions), the consensus time increases dramatically for low $k$-values. In this case, the stubborn agent effectively neutralizes the natural tendency of the population to converge to the 0-state.
IV Discussion
Our results demonstrate that opinion dynamics are significantly influenced by the asymmetry of interaction neighborhoods. In symmetric networks, consensus probabilities and times are unaffected by connectivity. However, in asymmetric networks, the probability of consensus depends strongly on $k$, and convergence times differ markedly between the two states.
Specifically, in the case of a randomly and equally distributed initial population of binary agents, the probability of reaching consensus behaves differently in symmetric and asymmetric scenarios. In the symmetric case, the probability of consensus for either state is equal (1/2) and remains independent of connectivity. In contrast, for the asymmetric case, the probability becomes an increasing function of connectivity, scaling with the population size (see Fig. 3).
Regarding the time required to achieve consensus, such as 0-consensus (where all agents adopt state 0), the behavior also differs. For the symmetric case, consensus time decreases with increasing network connectivity ($k$). However, in the asymmetric case, consensus time increases with kk, exhibiting a minimum when both states are considered equally likely to achieve consensus (see Fig. 4).
This asymmetry is particularly pronounced when a stubborn agent is included in the population. If the stubborn agent holds state 0 and interacts with its nearest neighbors, the consensus time is significantly reduced. Conversely, if the stubborn agent permanently holds state 1, with a neighborhood consisting of the friends of its closest friends (excluding the latter), the time to achieve consensus in state 1 increases dramatically for low $k$-values (see Fig. 5).
These results demonstrate that incorporating an asymmetric neighborhood significantly alters the opinion dynamics of the population, leading to several reflections on potential applications of this model, some of which are explored in the next section.
V Concluding remarks
The model and results presented in this paper could be particularly relevant in situations where contrasting opinions from different sources determine the final outcome. A familiar example, as described in the introduction, involves the advice provided by close friends compared to that of less familiar individuals. Close friends often tend to be more polite or hesitant to offer critical feedback, whereas friends-of-friends may provide more honest, albeit less personalized, opinions. This dynamic can sometimes lead to misleading decisions.
In a broader context, this model can help understand how misinformation—or highly polarized opinions—spreads through direct, peer-to-peer connections (first-degree neighbors), while fact-checking or corrective information may rely on more extended networks (second-degree neighbors or more distant sources). Misinformation tends to thrive within immediate social circles, reinforcing echo chambers, as individuals often prioritize interactions within their close-knit networks. Conversely, people seeking accurate information or corrections may need to rely on second-degree connections, such as fact-checkers or experts, who lie outside their immediate peer group. In this context, the model offers insights into the persistence of polarized views and explains why corrective information often takes longer to propagate through social networks.
On a similar note, the model could also provide valuable insights into political campaign strategies. It could be used to study how certain ideological groups primarily interact within their close communities (first-degree neighbors), while others may be more influenced by second-degree connections, such as political analysts or external media outlets. The findings from this study could guide campaigners in effectively targeting specific groups by adopting tailored outreach strategies. For example, campaigns could focus on leveraging first-degree influencers for tight-knit communities or second-degree influencers for more diffuse audiences, depending on the nature of the opinion being addressed.